OpenStudy (anonymous):

HERE is a challange(for dumb people like me) Let g(x)=log2(x) ;;;;; Find g^(-1) (3)

OpenStudy (anonymous):

Is that:\[g\left(x\right)=\log_2\left(x\right)\]? If it is, inverting the function...\[2^{g\left(x\right)}=x\]And evaluating at x=3

OpenStudy (anonymous):

that confuses me :/ **sorry** >_<

OpenStudy (anonymous):

Don't apologize for imagined issues. Anyways, \(g^{-1}\left(x\right)\) generally suggests the functional inverse, not \(\frac{1}{g\left(x\right)}\). So using the identity \(\log_a b=c\to a^c=b\)...